# Deep RL Dual Sourcing Inventory Management with Supply and Capacity Risk Awareness

### v3

@@ 2 Related work
Reinforcement Learning for Inventory Control: Recently, the progress of Reinforcement Learning (RL) on various optimization problems (Bello et al., 2016; Silver et al., 2016; Nazari et al., 2018; Liu et al., 2022; Liu, 2022; Liang et al., 2024; Shao et al., 2024) motivates its application in inventory management problems. In this context, Deep RL approaches have emerged more recently over its other Model-Based counterparts, due to improved computational scalability and generalization performance of Deep Neural Network (DNN) architectures. Furthermore, DRL methods have been shown to achieve performance gains over benchmarks for the multi-period newsvendor problem under fairly realistic assumptions on costs, prices, demand and stationarity (Balaji et al., 2019). It is worth highlighting that DRL for single sourcing problem has been comprehensively investigated with extensive empirical evidences at Amazon (Madeka et al., 2022).
@@ 4.2 Main Results and Analytics
[…]
@@ 5.1 Forecasting the Capacity Prices by a Neural Coordinatator
$\mathbf{\mathcal{P}_{3}}:$ $~\underset{\omega\in\Omega}{\min}~\mathbb{E}\Big{[}\sum_{t=0}^{T-1}(\sum_{i\in A}v^{i}I_{t}^{i}-K_{t})_{+}^{2}+||\lambda_{t}||+\mathcal{L}\left(\lambda_{t},H_{t}^{\lambda}\right)\Big{]},$ (18)
@@
$\mathcal{L}\left(\lambda_{t},H_{t}^{\lambda}\right)=\sum_{s=1}^{L}||\lambda_{t}-(\hat{\lambda}_{t-s}^{L})_{L-s}||^{2},$ (20)

### v5

@@ 2 Related work
Reinforcement Learning for Supply Chain Optimization Recently, the progress of applying Machine Learning in various optimization problems (Bello et al., 2016; Silver et al., 2016; Khalil et al., 2017; Nazari et al., 2018; Liu et al., 2021; Bengio et al., 2021; Liu et al., 2022; Liu, 2022; Sinclair et al., 2023; Liu et al., 2023; Shao et al., 2024) motivates its application in inventory management problems (Madeka et al., 2022; Qi et al., 2023; Liang et al., 2024a, b; Chen et al., 2025). In this context, Deep RL approaches have emerged more recently over its other Model-Based counterparts, due to improved computational scalability and generalization performance of Deep Neural Network (DNN) architectures. Furthermore, DRL methods have been shown to achieve performance gains over benchmarks for the multi-period newsvendor problem under fairly realistic assumptions on costs, prices, demand and stationarity (Balaji et al., 2019). It is worth highlighting that DRL for single sourcing problem has been comprehensively investigated with extensive empirical evidences at Amazon (Madeka et al., 2022).
@@ 4.2 Main Results and Analytics
From the results above, we can observe that overall dual sourcing strategies DualSrc-RL, TBS are favorable in terms of rewards. Furthermore, DualSrc-RL is most profitable in all the run scenarios, and has $4\%$ reward gains over TBS.
@@ 5.1 Forecasting the Capacity Prices by a Neural Coordinatator
[…]
@@
[…]